# LLM Stinger: Jailbreaking LLMs using RL fine-tuned LLMs

### v1

@@ Results
Research on adversarial attacks against large language models (LLMs) has made considerable progress, with early approaches focusing on manually crafted prompts and white-box techniques to induce harmful outputs. As LLMs have become more prevalent, the need for automated, scalable attack methods has grown, leading to the development of more advanced strategies. Recent work has introduced a variety of techniques, including gradient-based optimization of adversarial suffixes, role-playing scenarios, and genetic algorithms designed to exploit model vulnerabilities. In this section, we review the major contributions in the field, emphasizing both the breadth of these approaches and the increasing sophistication of adversarial methods that challenge the alignment of LLMs.
Zou et al. focus on finding a suffix that induces LLMs to produce objectionable content when attached to the end of queries. The attacks involve using a combination of greedy and gradient-based search techniques to optimize this suffix string, an approach that leads to transferability across models (Zou et al. 2023). Mehrotra et al. talk about Tree of Attacks with Pruning (TAP), an automated method for generating jailbreak prompts while requiring only black-box access to the target. TAP continuously refines the prompt using an attacker LLM (Vicuna-13B). TAP uses tree-of-thought reasoning and an evaluator LLM (GPT-4) to quickly search through a large space and send only likely jailbreak prompts to the target (Mehrotra et al. 2023). Tian et al. introduce Evil Geniuses (EG), a role-playing attack method that leverages Red-Blue exercises to improve the aggressiveness of the generated prompt. The queries are autonomously produced through GPT-3.5 using a template-based attack strategy (Tian et al. 2023). Liu et al. have developed AutoDAN, an attack method that generates jailbreak prompts through a hierarchical genetic algorithm. This role-playing attack method uses selection, crossover, and mutation to modify the prompt. Each iteration is evaluated using a fitness score that guides future modifications (Liu et al. 2023). Chao et al. introduce Prompt Automatic Iterative Refinement (PAIR), a fully automated method for jailbreak prompt discovery. PAIR uses Mixtral 8x7B Instruct as an attacker LLM to generate adversarial prompts using role-playing templates. A judgment LLM, Llama Guard in this case, provides a score, which is passed back into the attacker LLM to create an automated loop (Chao et al. 2023). Wang et al. present a thorough evaluation of the trustworthiness of recent GPT models, particularly GPT-4 and GPT-3.5, across various dimensions. It highlights that while GPT-4 generally outperforms GPT-3.5 in standard benchmarks, it is more susceptible to jailbreaking prompts (Wang et al. 2023). Wei et al. focus on evaluating the susceptibility of state-of-the-art language models to jailbreak attacks. These attacks target both suffix-based and role-playing techniques, using methods such as prefix injection, refusal suppression, Base64 encoding, and style injection. The paper argues that jailbreaks may be inherent in existing safety training methods and that scaling up models does not resolve these issues.  (Wei, Haghtalab, and Steinhardt 2024). Yao et al. introduce FuzzLLM, a universal fuzzing framework designed to proactively identify and evaluate jailbreak vulnerabilities in LLMs. By leveraging templates that combine jailbreak constraints with prohibited questions, FuzzLLM automates the testing process, significantly reducing manual effort and improving the efficiency and comprehensiveness of vulnerability discovery across various LLMs (Yao et al. 2024). Chu et al. perform a comprehensive assessment of various jailbreak attack methods on LLMs, revealing vulnerabilities bypassing safeguards. It evaluates 13 jailbreak methods across four categories, 160 questions from 16 violation categories, and six popular LLMs, finding that optimized prompts yield high attack success rates (Chu et al. 2024). Sun et al. introduce TrustLLM, a comprehensive study on the trustworthiness of LLMs, proposing principles across eight dimensions of trustworthiness and establishing a benchmark across six dimensions: truthfulness, safety, fairness, robustness, privacy, and machine ethics. It evaluates 16 mainstream LLMs using over 30 datasets, finding a positive correlation between trustworthiness and utility, with proprietary LLMs generally outperforming open-source ones in terms of trustworthiness (Sun et al. 2024). Zhou et al. introduce the InContext Adversarial Game (ICAG), a novel framework for defending Large Language Models (LLMs) against jailbreak attacks without requiring fine-tuning. By leveraging agent learning, ICAG employs an iterative adversarial game between attack and defense agents (GPT-3.5), dynamically refining defense mechanisms against evolving jailbreak prompts. This approach contrasts with static dataset-dependent methods, offering a more adaptable and effective defense strategy (Zhou et al. 2024). Xu et al. conduct a thorough analysis of jailbreak attacks on Large Language Models (LLMs) and their defense techniques. It evaluates nine attack methods and seven defense strategies across three LLMs: Vicuna, LLama, and GPT-3.5 Turbo. The study reveals that white-box attacks are less effective than universal techniques and that special tokens significantly influence the success of attacks.  (Xu et al. 2024). Geiping et al. explore a wide range of adversarial attacks on Large Language Models (LLMs), extending beyond alignment bypassing to include extraction, misdirection, denial of service, and control flow attacks. By employing optimization algorithms, the authors demonstrate how LLMs can be coerced into revealing system prompts, generating malicious URLs, and executing harmful instructions, despite safety training.  (Geiping et al. 2024). Pasquini et al. introduce Neural Exec, a new family of prompt injection attacks that autonomously generate execution triggers through an optimization-driven approach. These triggers are more effective and flexible than handcrafted ones and can bypass common preprocessing operations in Retrieval-Augmented Generation (RAG) pipelines. The paper demonstrates that these triggers can deviate significantly from known attacks, challenging existing blacklist-based defenses.  (Pasquini, Strohmeier, and Troncoso 2024).
@@
Previous jailbreak attacks have either focused on using white-box gradient-based techniques to find suffix strings, or prompt-level manipulations in conversational settings (such as role-playing templates), or manually figuring out certain jailbreak patterns. In most of these cases, the attacker model is not fine-tuned. In contrast, we are the first to fine-tune an attacker LLM to generate a new and modified version of existing suffix-based attacks. This feedback-driven approach enables the automated evolution of attacks, marking a new direction in adversarial methods for LLMs.
@@
The use of a language model in an RL loop to detect web vulnerabilities has been explored by BertRLFuzzer (Jha et al. 2023), a fuzzing tool that combines the BERT language model with RL. It uses binary feedback from an attack detector to automatically generate grammar-adhering mutations, producing attack vectors to uncover security vulnerabilities in web applications. In contrast, LLM Stinger leverages a large language model to generate complete attack vectors and utilizes fine-grained feedback signals, drawing inspiration from the Reinforcement Learning with Symbolic Feedback approach (Jha et al. 2024).
@@ Conclusions
We present LLM Stinger, a new adversarial attack framework designed to jailbreak safety-trained LLMs LLMs by generating adversarial suffixes within an RL loop. Tested against 15 attack methods across 7 LLMs, LLM Stinger showed a 52.2% attack success rate (ASR) on Claude 2, a model specifically designed to prevent such attacks, compared to the next best method’s 1.9% ASR. The system’s scalability and efficacy demonstrate its capability to break through various LLMs using only black-box API access.

### v2

@@ Results
Research on adversarial attacks against large language models (LLMs) has progressed rapidly, evolving from manually crafted prompts and white-box attacks to fully automated and scalable methods. As LLM deployment has expanded, so has the need to systematically understand, generate, and defend against jailbreak attacks. Prior work spans suffix-based optimization, role-playing and conversational attacks, evolutionary and fuzzing-based methods, large-scale evaluations, and emerging defense strategies. We organize the related work into thematic categories below.
Suffix-Based and Optimization-Driven Attacks
Zou et al. (2023) introduce universal adversarial suffixes optimized using greedy and gradient-based search techniques, demonstrating strong transferability across models. Geiping et al. (2024) generalize optimization-based attacks beyond alignment bypassing to include prompt extraction, misdirection, denial of service, and control flow manipulation, showing that safety training alone is insufficient to prevent coercive behaviors. Pasquini et al. (2024) propose Neural Exec, which uses optimization-driven prompt injection triggers that can bypass preprocessing stages in Retrieval-Augmented Generation pipelines and evade blacklist-based defenses.
Role-Playing and Conversational Jailbreak Attacks
Tian et al. (2023) propose Evil Geniuses, a Red-Blue team inspired role-playing framework that autonomously generates aggressive jailbreak prompts using GPT-3.5. Liu et al. (2023) introduce AutoDAN, which employs a hierarchical genetic algorithm with selection, crossover, and mutation operators to iteratively refine role-playing jailbreak prompts based on a fitness score. Chao et al. (2023) propose Prompt Automatic Iterative Refinement (PAIR), which uses an attacker LLM to generate role-playing prompts and a judgment LLM to provide feedback, forming a fully automated refinement loop.
Tree-Based, Evolutionary, and Fuzzing Approaches
Mehrotra et al. (2023) propose Tree of Attacks with Pruning (TAP), which uses tree-of-thought reasoning to iteratively refine jailbreak prompts under black-box access constraints, guided by an evaluator LLM. Yao et al. (2024) introduce FuzzLLM, a universal fuzzing framework that combines jailbreak constraints with prohibited queries to systematically identify vulnerabilities across multiple LLMs. Chu et al. (2024) provide a comprehensive empirical evaluation of 13 jailbreak methods across multiple models and violation categories, demonstrating that optimized and automated prompts consistently achieve high attack success rates.
Evaluation and Measurement of Jailbreak Vulnerabilities
Wang et al. (2023) analyze GPT-3.5 and GPT-4 across multiple trust dimensions, observing that stronger models may be more susceptible to jailbreaking despite improved benchmark performance. Wei et al. (2024) conduct a large-scale evaluation of suffix-based and role-playing jailbreaks, arguing that jailbreak vulnerabilities may be inherent to current safety training paradigms and are not resolved through model scaling alone. Sun et al. (2024) introduce TrustLLM, a benchmark spanning multiple trust dimensions and models, finding correlations between trustworthiness and utility. Xu et al. (2024) systematically compare attack and defense methods across several LLMs, highlighting the importance of universal attacks and the role of special tokens in jailbreak success.
Defense Strategies Against Jailbreak Attacks
Zhou et al. (2024) propose the InContext Adversarial Game (ICAG), which formulates defense as an iterative adversarial interaction between attacker and defender agents without requiring fine-tuning. This agent-based framework enables adaptive defenses against evolving jailbreak strategies.
Learning-Based Attack Generation and Reinforcement Learning
@@
Most prior jailbreak attacks either rely on white-box gradient-based optimization, prompt-level manipulations such as role-playing, or manually identified jailbreak patterns. In these approaches, the attacker model itself is not fine-tuned. In contrast, we are the first to fine-tune an attacker LLM to generate new and modified versions of suffix-based attacks. Our feedback-driven framework enables automated attack evolution, introducing a new paradigm for adversarial LLM research.
@@
The use of language models within reinforcement learning loops for vulnerability discovery has been explored in other domains. Jha et al. (2023) combine a BERT-based language model with reinforcement learning and binary feedback to generate grammar-adhering mutations for web vulnerability discovery. Our approach differs by leveraging a large language model to generate complete attack vectors and by using fine-grained symbolic feedback signals, drawing inspiration from Reinforcement Learning with Symbolic Feedback (Jha et al. 2024).
@@ Conclusions
We present LLM Stinger (Jha et al. 2025), a new adversarial attack framework designed to jailbreak safety-trained LLMs LLMs by generating adversarial suffixes within an RL loop. Tested against 15 attack methods across 7 LLMs, LLM Stinger showed a 52.2% attack success rate (ASR) on Claude 2, a model specifically designed to prevent such attacks, compared to the next best method’s 1.9% ASR. The system’s scalability and efficacy demonstrate its capability to break through various LLMs using only black-box API access.